The hypotenuse of right triangle is 15mc. One of the legs is 3cm more than the other. Find the length of the two legs.

Answer by amarjeeth123(569)

Let the two legs be x and x+3.
The hypotenuse of right triangle is 15cm.
For a right angled triangle the square of the hypotenuse is equal to the sum of squares of the other two sides.
Plugging in the values we get, x^2+(x+3)^2=15^2
Simplifying we get,
2x^2+6x+9=225
2x^2+6x+9-225=0
2x^2+6x-216=0
x^2+3x-108=0
x^2+12x-9x-108=0
x(x+12)-9(x+12)=0
(x-9)(x+12)=0
x-9=0 or x+12=0
x=9 or x=-12
x=9 is the correct answer as x cannot be negative.
The two legs of the triangle are 9 and 12 respectively.